OpenStudy (sighn0more):

Hello! Please help me work through this quadratic equation (it must be solved by completing the square): 2x^2 + 4x + 1 = 0

OpenStudy (anonymous):

do you have to complete the square here?

OpenStudy (sighn0more):

Yes! So far I have moved the constant to the LHS and have factored out the 2. I don't know where to go from there.

OpenStudy (anonymous):

don't factor out the 2 you have an equation, divide by 2

OpenStudy (anonymous):

\[2x^2 + 4x + 1 = 0\\2x^2 + 4x =-1\\ x^2+2x=-\frac{1}{2}\]

OpenStudy (anonymous):

then complete mr square half of 2 is 1, and 1 squared is 1 go to \[(x+1)^2=-\frac{1}{2}+1\]

OpenStudy (sighn0more):

Ah, I see now. Okay, so from there I would have (x + 1)^2 = 1/2 and then x = -1 + or - radical 1/2

OpenStudy (anonymous):

yes

OpenStudy (sighn0more):

Thank you! This helped a lot. Do you think that you can help me with one more problem? At least half way?